OpenStudy (wondermath):

Determine all real values of x given that 6^(x+1)-6^x=3^(x+4)-3^x

OpenStudy (anonymous):

apparently the answer is 4, but i am not sure exactly how. somehow this can be rewritten. let me try with paper

OpenStudy (anonymous):

\[6^x(6-1)=3^x(3^4-3)\] maybe that is a good start

OpenStudy (anonymous):

then \[5\times 6^x=80\times 3^x\]

OpenStudy (anonymous):

\[6^x=16\times 3^x\] \[\frac{6^x}{3^x}=16\] \[2^x=16\] \[x=4\] there i knew it was possible

OpenStudy (wondermath):

thanks!